I am solving this problem:
Suppose that you take all of the black cards out of a standard deck of 52 cards and throughly shuffle the remaining 26 red cards. From this deck of 26 red cards you will select 2 cards, one at a time, without replacement, and record whether each card picked is a face card (king, queen, jack) or not a face card.'
I would like to know: what is the probability that none of the two cards picked in this way is a face card.
I know the equation is something like 20C2/26C2, but i do not know what "C2" stands for.
Thank you for any help you can provide!

Answer by stanbon(75887)

I know the equation is something like 20C2/26C2, but i do not know what "C2" stands for.
-----
20C2 is the number of pairs you get from 20 distinct objects.
26C2 is similar.
----------------
20C2 = (20*19)/(1*2) = 190
26C2 = (26*25)/(1*2) = 13*25
----------
Cheers,
Stan H.
------